Planning week 9 - 16 Jun 2018

This planning week took place during the Hinode eclipse season. It began with no activity, and only a very weak, decaying AR at central meridian, which was the default target initially. On 11 June, AR 12713 appeared over the east limb, and this became the default target although it produced no significant flaring activity.

There were a few irritations this week:

  • XRT is running a new disk center synoptic program between 0 and 3 UT every day.
  • eis_mk_plan is very slow when saving studies on the timeline (seems only to affect computers with MacOS 10.13.5?). This was a particular problem for HOPs 79 and 130.
  • HOP 358 (with GREGOR and SST) only sent pointing information in time for the final plan, with no indication of what they'd be looking at.
  • SOT was running long calibration studies at disk center.

Interesting studies (for me) were:

  • Isolated fan loop study on 9 & 10 June.
  • North coronal hole map on 16 June.
  • Tracking of AR with 8 HPW021 rasters on 14 June.

Saturday 9 June (OP1)

AIA 193 image

HOP 358 (GREGOR & SST) was observing the south coronal hole, so I scheduled an Atlas_60 at 11:27 UT.

There was bright fan loop in the FOV, so I scheduled cool_loop_response four times at 18:23, 19:47, 21:15 and 22:54 UT. There was no SOT or XRT data. UPDATE: this loop survived and the data should be good.

Sunday 10 June (OP1)

AIA 193 image

HOP 333 (center-to-limb quiet Sun with ALMA) was partially supported between 15:20 and 17:00. The PI gave four pointings but due to eclipses I could only support the first three using the study an_alma_eis_s01v01.

I continued the fan loop study with cool_loop_response rasters at 19:02, 20:12 and 21:51 UT. UPDATE: the fan was just about gone as the first raster started.

Monday 11 June (OP1)

AIA 193 image

HOP 336 (cycle 25 bright points) was supported at two pointing locations between 12:37 and 19:00. I ran Large_CH_Map twice at each location.

I ran two rasters of HPW021VEL260x512v2 at 20:57 and 22:33 UT at the east limb active region.

Tuesday 12 June (OP1)

AIA 193 image

A quiet Sun atlas was scheduled at 00:04 UT.

Two Atlas_60's were performed for the AR at the east limb at 03:21 and 06:38 UT. The first was partly on disk, and the second was set about 100" further out.

Tuesday 12 June (OP2)

AIA 193 image

HOP 130 was scheduled from 12:00 to 19:00.

A HPW021 raster was run on the east limb AR at 21:31 UT

Five orbits were used for quiet Sun disk center tracking between 23:08 and 06:40 using dhb_atlas_120m_30.

Wednesday 13 June (OP2)

AIA 193 image

The quiet Sun tracking continued until 06:40.

For the AR at the east limb, I used cool_loop_response between 19:40 and 24:00. I'm hoping to catch the footpoints at the leading edge of the AR.

Thursday 14 June (OP2)

AIA 193 image

A second HPW021 raster of the OP was run on the east limb AR at 02:53 UT.

Thursday 14 June (OP3)

AIA 193 image

EIS had 30% of the DR for this OP. I scheduled 4 then 3 runs of HPW021 in 7 of 8 orbits beginning at 11:07 (sequence was interrupted by a XRT synoptic).

Friday 15 June (OP3)

AIA 193 image

HOP 79 began at 03:30 with 20 pointings (N-S) and continued until 02:30 on Saturday.

Saturday 16 June (OP3)

AIA 193 image

I took advantage of a SOT QS pointing to take two Large_CH_map studies at the north pole. The first at 02:32 is set to cover the upper part of the FOV (including off-limb), and the second at 04:04 UT covers the lower part, including a CH extension.

I scheduled two more HPW021 rasters on AR 12713 at 05:44 and 08:59 UT.

Page maintained by Dr Peter R Young.